Tony has twice as much money as Alicia. She has $16 less than Ralph. Together they have $200. How much money does each have?
Answer by drj(1380)

Tony has twice as much money as Alicia. She has $16 less than Ralph. Together they have $200. How much money does each have?

Step 1. Let x=amount of dollars that Alicia has.

Step 2. Tony has twice as much as Alicia or 2x dollars

Step 3. Ralph has x+16 dollars since Alicia has $16 less than Ralph.

Step 4. Since the total amount is $200 among Alicia, Tony and Ralph, add Steps 1-3. x+2x+x+16=200. This yields. 4x+16=200 or 4x=184 where I subtracted 16 from both sides of the equation.

Step 5. Solving x=184/4=46 dollars.

a. Alicia has 46 dollars
b. Tony has 46*2=92 dollars.
c Ralph has 46+16=62 dollars.
d. Total should be $200. So, 46+92+62=200. So it checks out.

For Step-By-Step Videos in Introductory Algebra and more word problems, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra.